Consider the model y = a + bX + e. Show that the least squares estimator for b is unbiased and consistent. You can assume that the 5 standard disturbance term assumptions are true. For each step explain why it is true.
Answer #1

Please give me a thumbs-up, if you are satisfied with my answer.
we have the model & Y= at bxte. of bis givers by the least square estimate 6- cxi-2) (Yi-y) according to the model, Y - atbxt thes (6) - E[b+ F(xi-T)ei 1 {(xi-xy +E[6] + E ( 263-) ei bis constant ( -X of linear model, xs are finced in they are not Byconsistency of last square estimator we have 6 = b + (xi-x) ei = 6 + 16 =(xi-x) ei In {(xi-x)2 multiply and divide by / in th